# MADNESS Deblender

### v1

@@ Abstract
Context. Due to the unprecedented depth of the upcoming ground-based Legacy Survey of Space and Time (LSST) at the Vera C. Rubin Observatory, approximately two-thirds of the galaxies are likely to be affected by blending — the overlap of physically separated galaxies in images. Thus, extracting reliable shapes and photometry from individual objects will be limited by our ability to correct blending and control any residual systematic effect. Deblending algorithms tackle this issue by reconstructing the isolated components from a blended scene, but the most commonly used algorithms often fail to model complex realistic galaxy morphologies.
@@
Methods. MADNESS is based on deep neural network architectures such as variational auto-encoders and normalizing flows. The variational auto-encoder reduces the high-dimensional pixel space into a lower-dimensional space, while the normalizing flow models a data-driven prior in this latent space. Together, these neural networks enable us to obtain the maximum a posteriori solution in the latent space.
@@ 1 Introduction
One of the main goals of the next generation cosmological surveys such as the Legacy Survey of Space and Time (LSST [^1] https://www.lsst.org/, LSST Science Collaboration et al., 2009) at the Vera C. Rubin Observatory, ESA’s Euclid [^2] https://www.esa.int/Science_Exploration/Space_Science/Euclid telescope (Refregier et al., 2010), and the Nancy Grace Roman Space Telescope [^3] https://roman.gsfc.nasa.gov/ (Spergel et al., 2013), is to understand the fundamental nature of dark energy. Among the cosmological probes that can be used for these studies, the use of cosmic shear, in particular, is expected to make a major step forward with these Stage IV surveys. Cosmic shear corresponds to the systematic distortions in the shapes of astrophysical objects due to the bending of the path of light by inhomogeneous matter distribution present between us and the distant source galaxies (so-called gravitational lensing). Studying these correlated distortions in tomographic redshift bins allows us to map the underlying dark matter distribution and its evolution in time, and thus constrain dark energy parameters. However, in the weak regime of gravitational lensing, these distortions imprinted on galaxies are extremely small, of the order of a few percent. This requires high statistical precision to impose strong constraints on dark energy and forces us to probe the universe to an unprecedented depth with these wide-field surveys.
@@
With increased statistical precision, we also expect new challenges in data analysis. One such challenge is the blending of galaxies which occurs when physically separated objects lie along the same line of sight and overlap in the pixels. The increase in object density due to greater depth of observations will make the effect of blending significant in upcoming surveys. In the case of ground-based surveys such as the LSST, the point-spread-function (PSF) contribution from the Earth’s atmosphere will lead to an additional blurring of images and increase the effect of blending. In the wide survey of Hyper Suprime-Cam (HSC [^4] https://hsc.mtk.nao.ac.jp/ssp/, Aihara et al., 2017) with a limiting i-band magnitude $\approx 26$, around $58\%$ of the objects were observed to be blended (Bosch et al., 2017). The problem is expected to be more severe in the case of LSST since it aims to reach $i\approx 27$ after 10 years of operation. As a consequence, it is expected that around $62\%$ of LSST objects will have at least $1\%$ flux contribution from neighboring sources (see Sanchez et al., 2021).
@@
The extent to which we can control the systematic errors will determine how well we can constrain cosmological models (see Mandelbaum, 2018, for a detailed discussion on weak lensing systematic effects), and blending is going to be a major part of these systematic errors. For example, blending can bias photometric redshift (photo-z) measurements that will in turn affect the downstream analysis of most cosmological probes including weak lensing. To measure galaxy properties on blended scenes, the effects of blending must be mitigated before processing the data with pipelines that are designed to operate with single sources. The inverse problem of reconstructing the individual isolated components from a blended scene is called deblending. Another approach for correcting weak lensing measurements is to calibrate for the effect of blending, among other systematic effects, using image simulations (MacCrann et al., 2021; Sheldon et al., 2020). However, the photometry of the sources is still affected and it remains to be demonstrated that these simulations would be sufficient to capture the entirety of the blending impacts for LSST weak lensing analysis.
@@
To make matters worse, a considerable fraction of these blends are likely to go undetected. These cases in which the detection algorithm fails to identify multiple sources among blended galaxies are called ‘unrecognized blends’. Typically, this can happen when the centers of two objects are closer to each other than a fraction of the PSF or when one of the sources is much brighter than the other so that the detection algorithm is not able to distinguish the fainter source. While it is more optimal to perform joint detection and deblending, combining the two steps can be computationally very expensive. Most deblenders assume prior detection of objects; therefore they rely on robust detection algorithms to deblend galaxies from a blended scene efficiently. Synergies between different surveys can also mitigate the effects of unrecognized blends. Since space-based surveys have no atmospheric contribution to the PSF the sources are more resolved and can help provide priors for detections in ground-based surveys where the blending effect is stronger due to the atmosphere (see Melchior et al., 2021). In fact, Arcelin et al. (2020) showed that joint pixel-level processing of ground and space-based data can help improve the quality of deblended models, motivating us to further develop algorithms that are capable of simultaneously processing data from multiple surveys.
@@
One of the most commonly used deblending algorithms, called SExtractor and developed by Bertin & Arnouts (1996), performs image segmentation by looking for connected components at various thresholds and computes the deblending solution independently on each band or uses a dual-mode to use one band as the reference (e.g., the r-band) to obtain photometry in another. Such a method cannot realistically deblend a galaxy because it will assign pixels in an overlapping case to one of the galaxies without trying to redistribute the fluxes. This results in strong limitations in terms of both morphology and fluxes of the reconstructions.
@@
Although the Sloan Digital Sky Survey (SDSS) deblender by Lupton (2005) took a step further and assumed symmetry properties of galaxies as priors, it still performed the deblending individually in different bands. Other approaches like Lambda Adaptive Multi-Band Deblending Algorithm in R (LAMBDAR, Wright et al., 2016) use higher-resolution aperture priors to obtain consistent matched photometry across a range of photometric data with various resolutions. Another commonly used algorithm called TheTractor (Lang et al., 2016) optimizes over models of astronomical objects to find the optimal model that maximizes the likelihood. More recently, Melchior et al. (2018) have developed the scarlet deblender that combines information from different bands and uses a constrained matrix factorization to separate the overlapping sources. scarlet uses proximal gradient descent to impose constraints such as symmetry and monotonicity of light profiles on the reconstructed components. The scarlet deblender makes it possible to incorporate physical information such as noise and the PSF into the algorithm so that we can leverage knowledge about the data. While such an approach is capable of handling more complex blended scenes and provides a significant improvement over previous deblenders, modeling complex morphologies is still a challenge because of the strong regularizations.
@@
To predict more complex galaxy shapes, the use of deep learning is gaining ground. In an approach very similar to scarlet and also closely related to our work, Lanusse et al. (2019) replaced the regularizations by modeling a data-driven prior using differentiable deep generative models with explicit likelihoods. This approach solves the maximum-a-posteriori (MAP) solution of the inverse problem by using autoregressive models such as the PixelCNN++ (Salimans et al., 2017) to compute a likelihood based on a pixel-level galaxy prior. Although such a deep-learning architecture is quite interpretable, the pixel-wise auto-regressive nature negatively impacts the speed which is a major concern for upcoming surveys such as the LSST.
@@
Several other deep learning architectures based on generative models have already shown promising deblending results. Reiman & Göhre (2019) used a branched Generative Adversarial Network (GAN) to deblend two galaxies from a stamp. Arcelin et al. (2020) used a modified variational auto-encoder (VAE) architecture in which the encoder is trained to remove the non-central galaxies and project only the central galaxy into the latent space; the decoder is trained to map from the latent space representation of an isolated galaxy to its image. Using a similar VAE model as one of its components, Hansen et al. (2022) proposed Bayesian Light Source Separator (BLISS), which is a scalable approach to perform Bayesian detection and deblending using a combination of three convolutional encoder networks responsible for detection, star/galaxy classification, and morphology estimation. Unlike other approaches, BLISS aims to produce a probabilistic catalog of galaxies that can be used by downstream scientific analysis. More recently, Sampson et al. (2024) introduced scarlet2 which uses neutral score matching to obtain the priors for scarlet in order to predict more complex morphologies. Going beyond generative models, other deep learning techniques have also been shown to be effective for deblending. For example, Wang et al. (2022) uses a residual dense neural network to extract galaxies iteratively from a field until no galaxies are found.
@@
Deep learning methods using CNNs give us an easy way to combine multichannel and multi-instrument information (Arcelin et al., 2020), but although architectures that deblend with a single forward pass are extremely fast at obtaining results, they lack explicit optimization over the total flux. To find a middle ground between the VAE deblender proposed by Arcelin et al. (2020) and MAP estimate obtained by Melchior et al. (2018) and Lanusse et al. (2019), we have developed a deblender called MADNESS, which stands for Maximum A posteriori solution with Deep NEural networks for Source Separation. Although our work can be extended to obtain a probabilistic catalog and potentially be combined with the detection and classification steps of BLISS, the focus, for now, is on obtaining the MAP solution. Like most deblenders described above, MADNESS assumes that the galaxies have already been detected and the goal is to reconstruct the isolated components. As an extension to the deblender proposed by Arcelin et al. (2020), we couple the VAE with a normalizing flow (NF) introduced by Rezende & Mohamed (2015). The combination of the two neural networks (NN) is used in a way such that the MAP estimate can be obtained by a gradient descent in the latent space. By doing so, we show that deep learning can be used efficiently to perform an optimization rather than a single pass through a feed-forward network.
@@
We first describe the simulations used for training and testing MADNESS in Section 2. In Section 3 we dive deeper into variational autoencoders and normalizing flows and discuss in detail the architecture, training of the models, and the methodology of our deblender. We present the results of our algorithm and compare it with state-of-the-art deblenders in Section 4 using various metrics. Finally, in Section 5 we discuss potential next steps to prepare our algorithm for processing real data, before concluding in Section 6.
@@ 2 Dataset
Figure 1: Distribution of galaxies from the CatSim catalog before (orange) and after (blue) the SNR cut.
(a) Galaxies from IsolatedSims dataset.
(b) Galaxies from BlendedSims dataset.
Figure 2: Examples of normalized simulations in the r-band, used for training and validating the variational autoencoder and the normalizing flow. The central galaxy in the stamps is marked in black while the neighbors in the BlendedSims are marked in red. Each row represents the same scene and the first column shows the noisy field; the second column shows the isolated noiseless central galaxy; the third column shows the residual field after subtracting the central galaxy from the noisy field. The color bar for each plot shows the normalized electron counts in the pixels.
To train and test our deblender, we use galaxies from the CatSim  [^5] https://www.lsst.org/scientists/simulations/catsim galaxy catalog that has been prepared for the LSST catalog simulator (Connolly et al., 2014). The galaxies in the catalog are defined as a sum of bulge and disk models covering a redshift range of $60<z<$ 6 with a realistic distribution of shapes and sizes. All parameters of the galaxies are assumed to be the same in all bands except the magnitudes. For this study, we use a one-square-degree subset that has been previously used by the community to understand the effects of blending on cosmic shear estimation (Sanchez et al., 2021). This one square degree of data consists of $\approx 858k$ galaxies with a magnitude up to $28$ in the r-filter. We further impose a signal-to-noise ratio (SNR) cut of 10 on the r-filter computed using the SNR definition described in Appendix A. Figure 1 shows the distribution of redshift and r-AB magnitude of the galaxies before and after the SNR cut. Finally, we are left with roughly $317k$ galaxies of which we keep 200 000200000200\,000200 000 for training and validation and the remaining for testing.
LSST is expected to observe the sky with two $15-$second exposures, each per visit. The galaxies are simulated under $10$-year LSST conditions with a gain of $1$, standard zero-point, and background sky brightness computed using the expected number of visits in each filter of LSST (Željko Ivezić et al., 2019). The PSF in each band is considered constant over all the generated images but is different for different bands; the width of the PSF is taken from Željko Ivezić et al. (2019), assuming an airmass of $1.2$. These conditions for the simulations are fetched from SurveyCodex  [^6] https://github.com/LSSTDESC/surveycodex and summarized in table 1.
@@
Filter zero point sky brightness PSF FWHM visits
@@ 2.1 Training and validation data
We use the Blending-Toolkit  [^7] https://github.com/LSSTDESC/BlendingToolKit/ (BTK, Mendoza et al. in prep) to generate postage stamps for LSST as it allows us to easily control the number and location of galaxies in the stamps. BTK uses GalSim  [^8] https://github.com/GalSim-developers/GalSim (Rowe et al., 2015) internally to simulate the galaxies. Using BTK we generate stamps of size $45\times 45$ pixels in each band, which corresponds to 9 ″$\times$ 9 ″, and the galaxy to be deblended is always located at the center.
Of the 200 000200000200\,000200 000 galaxies kept for training and validation, we use $75\%$ for training and the remaining $25\%$ for validation. We generate two distinct datasets: IsolatedSims which contains only one galaxy located at the center of each stamp; BlendedSims consists of stamps that are extracted from a field with up to 3 galaxies, the centers of which lie within a box of $20\times 20$ pixels. With each galaxy in the fields as a center, a stamp of $45\times 45$ pixels is extracted.
The purpose of the BlendedSims dataset is to retrain the encoder so that it can initialize the latent space for faster convergence of the MAP optimization, as described in Section 3.4. The exact distribution of the number of galaxies in the blends will not have a significant impact on the results of MADNESS. So we let the number of galaxies in the fields used for extracting BlendedSims vary uniformly between 1 and 3. In Figure 2 we show sample images used to train our network.

### v2

@@ Abstract
Context. Due to the unprecedented depth of the upcoming ground-based Legacy Survey of Space and Time (LSST) at the Vera C. Rubin Observatory, approximately two-thirds of the galaxies are likely to be affected by blending - the overlap of physically separated galaxies in images. Thus, extracting reliable shapes and photometry from individual objects will be limited by our ability to correct blending and control any residual systematic effect. Deblending algorithms tackle this issue by reconstructing the isolated components from a blended scene, but the most commonly used algorithms often fail to model complex, realistic galaxy morphologies.
@@
Methods. MADNESS is based on deep neural network architectures, namely variational auto-encoders and normalizing flows. The variational auto-encoder reduces the high-dimensional pixel space into a lower-dimensional space, while the normalizing flow models a data-driven prior in this latent space. Together, these neural networks enable one to obtain the maximum a posteriori solution in the latent space.
@@ 1 Introduction
One of the main goals of the next-generation cosmological surveys such as the Legacy Survey of Space and Time (LSST [^1] https://www.lsst.org/, LSST Science Collaboration et al., 2009) at the Vera C. Rubin Observatory, ESA’s Euclid [^2] https://www.esa.int/Science_Exploration/Space_Science/Euclid telescope (Refregier et al., 2010), and the Nancy Grace Roman Space Telescope [^3] https://roman.gsfc.nasa.gov/ (Spergel et al., 2013), is to understand the fundamental nature of dark energy. Among the cosmological probes that can be used for these studies, the use of cosmic shear, in particular, is expected to make a major step forward with these stage IV surveys. Cosmic shear corresponds to the systematic distortions in the shapes of astrophysical objects due to the bending of the path of light by inhomogeneous matter distribution present between us and the distant source galaxies (so-called gravitational lensing). Studying these correlated distortions in tomographic redshift bins allows one to map the underlying dark matter distribution and its evolution in time, and thus constrain dark energy parameters. However, in the weak regime of gravitational lensing, these distortions imprinted on galaxies are extremely small, of the order of a few percent. This requires high statistical precision to impose strong constraints on dark energy and forces the community to probe the universe to an unprecedented depth with these wide-field surveys.
@@
With increased statistical precision, new challenges in data analysis are also expected. One such challenge is the blending of galaxies, which occurs when physically separated objects lie along the same line of sight and overlap in the pixels. The increase in object density due to the greater depth of observations will make the effect of blending significant in upcoming surveys. In the case of ground-based surveys, such as the LSST, the point-spread-function (PSF) contribution from the Earth’s atmosphere will lead to an additional blurring of images and increase the effect of blending. In the wide survey of Hyper Suprime-Cam (HSC [^4] https://hsc.mtk.nao.ac.jp/ssp/, Aihara et al., 2017), with a limiting i-band magnitude $\approx 26$, around $58\%$ of the objects were observed to be blended (Bosch et al., 2017). The problem is expected to be more severe in the case of LSST since it aims to reach $i\approx 27$ after 10 years of operation. As a consequence, it is expected that around $62\%$ of LSST objects will have at least a $1\%$ flux contribution from neighboring sources (see Sanchez et al., 2021).
@@
The extent to which the systematic errors can be controlled will determine how well cosmological models can be constrained (see Mandelbaum, 2018, for a detailed discussion on weak lensing systematic effects), and blending is going to be a major part of these systematic errors. For example, blending can bias photometric redshift (photo-z) measurements that will in turn affect the downstream analysis of most cosmological probes, including weak lensing. To measure galaxy properties on blended scenes, the effects of blending must be mitigated before processing the data with pipelines that are designed to operate with single sources. The inverse problem of reconstructing the individual isolated components from a blended scene is called deblending. Another approach for correcting weak lensing measurements is to calibrate for the effect of blending, among other systematic effects, using image simulations (MacCrann et al., 2021; Sheldon et al., 2020). However, the photometry of the sources is still affected, and it remains to be demonstrated that these simulations would be sufficient to capture the entirety of the blending impacts for LSST weak lensing analysis.
@@
To make matters worse, a considerable fraction of these blends are likely to go undetected. These cases in which the detection algorithm fails to identify multiple sources among the blended galaxies are called “unrecognized blends”. Typically, this can happen when the centers of two objects are closer to each other than a fraction of the PSF or when one of the sources is much brighter than the other, so that the detection algorithm cannot distinguish the fainter source. Although it is more optimal to perform joint detection and deblending, combining the two steps can be computationally very expensive. Most deblenders assume prior detection of objects; therefore, they rely on robust detection algorithms to deblend galaxies from a blended scene efficiently. Synergies between different surveys can also mitigate the effects of unrecognized blends. Since space-based surveys do not have an atmospheric contribution to the PSF, the sources are more resolved and can help provide priors for detections in ground-based surveys, where the blending effect is stronger due to the atmosphere (see Melchior et al., 2021). In fact, Arcelin et al. (2020) showed that joint pixel-level processing of ground- and space-based data can help improve the quality of deblended models, motivating further development of algorithms that are capable of simultaneously processing data from multiple surveys.
@@
One of the most commonly used deblending algorithms, SExtractor (developed by Bertin & Arnouts, 1996), performs image segmentation by looking for connected components at various thresholds and computes the deblending solution independently on each band or uses a dual mode to use one band as the reference (e.g., the r-band) to obtain photometry in another. Such a method cannot realistically deblend a galaxy because it will assign pixels in an overlapping case to one of the galaxies without trying to redistribute the fluxes. This results in strong limitations in terms of both the morphology and the fluxes of the reconstructions.
@@
Although the Sloan Digital Sky Survey (SDSS) deblender by Lupton (2005) took a step further and assumed symmetry properties of galaxies as prior, it still performed the deblending individually in different bands. Other approaches, such as Lambda Adaptive Multi-Band Deblending Algorithm in R (LAMBDAR; Wright et al., 2016) use higher-resolution aperture priors to obtain consistent matched photometry across a range of photometric data with various resolutions. Another commonly used algorithm, called TheTractor (Lang et al., 2016), optimizes astronomical object models to find the optimal model that maximizes the likelihood. More recently, Melchior et al. (2018) have developed the scarlet deblender, which combines information from different bands and uses a constrained matrix factorization to separate overlapping sources. scarlet uses proximal gradient descent to impose constraints such as the symmetry and monotonicity of light profiles on the reconstructed components. The scarlet deblender makes it possible to incorporate physical information such as noise and the PSF into the algorithm so that knowledge about the data can be leveraged. Although such an approach is capable of handling more complex blended scenes and provides a significant improvement over previous deblenders, modeling complex morphologies is still a challenge because of the strong regularizations.
@@
The use of deep learning is gaining ground in the interest of predicting more complex galaxy shapes. In an approach very similar to scarlet and also closely related to our work, Lanusse et al. (2019) replaced the regularizations by modeling a data-driven prior using differentiable deep generative models with explicit likelihoods. This approach solves the maximum a posteriori (MAP) solution of the inverse problem by using autoregressive models such as the PixelCNN++ (Salimans et al., 2017) to compute a likelihood based on a pixel-level galaxy prior. Although such a deep learning architecture is quite interpretable, the pixel-wise auto-regressive nature negatively impacts the speed, which is a major concern for upcoming surveys such as the LSST.
@@
Several other deep learning architectures based on generative models have already shown promising deblending results. Reiman & Göhre (2019) used a branched generative adversarial network (GAN) to deblend two galaxies from a stamp. Arcelin et al. (2020) used a modified variational autoencoder (VAE) architecture in which the encoder is trained to remove non-central galaxies and project only the central galaxy into the latent space; the decoder is trained to map from the latent space representation of an isolated galaxy to its image. Using a similar VAE model as one of its components, Hansen et al. (2022) proposed Bayesian Light Source Separator (BLISS), which is a scalable approach to perform Bayesian detection and deblending using a combination of three convolutional encoder networks responsible for detection, star-galaxy classification, and morphology estimation. Unlike other approaches, BLISS aims to produce a probabilistic catalog of galaxies that can be used by downstream scientific analysis. More recently, Sampson et al. (2024) introduced scarlet2, which uses neutral score matching to obtain the priors for scarlet in order to predict more complex morphologies. Going beyond generative models, other deep learning techniques have also been shown to be effective for deblending. For example, Wang et al. (2022) used a residual dense neural network to extract galaxies iteratively from a field until no galaxies were found.
@@
Deep learning methods using convolutional neural networks provide an easy way to combine multichannel and multi-instrument information (Arcelin et al., 2020). Although architectures that deblend with a single forward pass are extremely fast at obtaining results, they lack explicit optimization over the total flux. To find a middle ground between the VAE deblender proposed by Arcelin et al. (2020) and MAP estimate obtained by Melchior et al. (2018) and Lanusse et al. (2019), we have developed a deblender called MADNESS  [^5] https://github.com/LSSTDESC/madness, release v1.0.0, which stands for Maximum A posteriori solution with Deep NEural networks for Source Separation. Although our work can be extended to obtain a probabilistic catalog and potentially be combined with the detection and classification steps of BLISS, the focus for now is on obtaining the MAP solution. Similar to most deblenders described above, MADNESS assumes that the central pixels of galaxies in the field are known and the goal is to reconstruct the isolated components. As an extension to the deblender proposed by Arcelin et al. (2020), we coupled the VAE with a normalizing flow introduced by Rezende & Mohamed (2015). The combination of the two neural networks is used in a way that the MAP estimate can be obtained by a gradient descent in the latent space. By doing so, we show that deep learning can be used efficiently to perform an optimization rather than a single pass through a feed-forward network.
@@
We first describe the simulations used for training and testing MADNESS in Section 2. In Section 3, we dive deeper into variational autoencoders and normalizing flows and discuss in detail the architecture, training of the models, and the methodology of our deblender. We present the results of our algorithm and compare it with state-of-the-art deblenders in Section 4 using various metrics. Finally, in Section 5 we discuss potential next steps to prepare our algorithm for processing real data, before concluding in Section 6.
@@ 2 Dataset
Figure 1: Distribution of galaxies from the CatSim catalog before (orange) and after (blue) the S/N cut for r-band S/N (see Appendix A), redshift (from catalog), and r-AB magnitude (from catalog).
Figure 2: Examples of normalized simulations in the r-band, used for training and validating the variational autoencoder and the normalizing flow. Panel (a) shows galaxies from the IsolatedSims dataset, while panel (b) shows galaxies from the BlendedSims dataset. The central galaxy in the stamps is marked in black while the neighbors in the BlendedSims are marked in red. Each row represents the same scene, and the first column shows the noisy field; the second column shows the isolated noiseless central galaxy; the third column shows the residual field after subtracting the central galaxy from the noisy field. The color bar for each plot shows the normalized electron counts in the pixels.
To train and test our deblender, we use galaxies from the CatSim  [^6] https://www.lsst.org/scientists/simulations/catsim galaxy catalog that has been prepared for the LSST catalog simulator (Connolly et al., 2014). The galaxies in the catalog are defined as a sum of bulge and disk models covering a redshift range of $0<z<6$ with a realistic distribution of shapes and sizes. For each galaxy, all parameters related to shape and size are assumed to be the same in all bands, but not the magnitudes. For this study, we used a one-square-degree subset that has been previously used by the community to understand the effects of blending on cosmic shear estimation (Sanchez et al., 2021). This one square degree of data consists of $\approx 858k$ galaxies with a magnitude up to $28$ in the r-filter. We further imposed a signal-to-noise ratio (S/N) cut of 10 on the r-filter using the S/N definition described in Appendix A. Figure 1 shows the distribution of redshift and r-AB magnitude of the galaxies before and after the S/N cut. Finally, we are left with roughly $317k$ galaxies of which we keep 200 000200\,000 for training and validation and the remaining for testing.
The galaxies are simulated under $10$-year LSST conditions with two $15$-second exposures, each per visit, a gain of $1$, standard zero-point, and sky background brightness computed using the expected number of visits in each filter of LSST (Ivezić et al., 2019). Besides the sky background, we also added Poisson noise due to the flux in each pixel. We used the Kolmogorov model for the PSF, which is held constant over all the generated images but is different for each band; the width of the PSF is taken from Ivezić et al. (2019), assuming an airmass of $1.2$. These conditions for the simulations are fetched from SurveyCodex  [^7] https://github.com/LSSTDESC/surveycodex Version 1.2.0 and summarized in Table 1.
@@
Filter zero point sky background PSF FWHM visits
@@ 2.1 Training and validation data
We used the Blending-Toolkit  [^8] https://github.com/LSSTDESC/BlendingToolKit/ (BTK, Mendoza et al., 2025) to generate postage stamps for LSST as it allows easy control over the number and location of galaxies in the stamps. BTK uses GalSim  [^9] https://github.com/GalSim-developers/GalSim (Rowe et al., 2015) internally to simulate the galaxies. Using BTK we generated stamps of size $45\times 45$ pixels in each band, which corresponds to 9 $\times$ 9 ″, and the center of the galaxy to be deblended is always located within the central pixel of the stamp.
Of the 200 000200\,000 galaxies kept for training and validation, we used $75\%$ for training and the remaining $25\%$ for validation. We generated two distinct datasets: IsolatedSims, which contains only one galaxy located at the center of each stamp; BlendedSims, consisting of stamps that are extracted from a field with up to 3 galaxies, the centers of which lie within a box of $20\times 20$ pixels. With each galaxy in the fields as a center, a stamp of $45\times 45$ pixels is extracted.
The purpose of the BlendedSims dataset is to retrain the encoder so that it can initialize the latent space for faster convergence of the MAP optimization, as described in Section 3.4. The exact distribution of the number of galaxies in the blends will not have a significant impact on the results of MADNESS. So we let the number of galaxies in the fields used for extracting BlendedSims vary uniformly between 1 and 3. In Figure 2, we show sample images used to train our network.